# Review of performance metrics of spin qubits in gated semiconducting nanostructures

### v7

@@ I.2 Spin-qubit types, device geometries, material choices
Every value given in this Technical Review (such as a coherence time of 10 ns) is a result of a measurement with a specific device and qubit type, reported in a published reference. [^1] Note that whereas we count arXiv preprints as ‘published references’, the majority of the entries are from peer-reviewed journal publications. The current bibliography for the database contains 13 arXiv preprints, only four of which were uploaded before the year 2023, Refs. [120, 46, 286, 99]. This means, first of all, that we include only numbers explicitly stated and directly measured. We do not include extrapolations to other conditions or materials. Also, we do not usually derive the values ourselves even if it would be possible. For example, if the reference gives the operation time and the dephasing time, but does not state the quality factor, we do not evaluate the latter ourselves. However, if a quantity is discussed and presented as a figure (for example, the operation time implied from oscillations displayed by a resonantly driven qubit), we might include a value read-off from the figure. In such cases, the table entry contains a ‘Note’ with keywords (derived)  or (estimated) , which are explained in Appendix D. With this requirement, every value given in this Technical Review should be easy to find using the reference, given under the attribute ‘Reference’, and within the reference as described by the attribute ‘Source’. An example of the latter is say “page 4 and Fig. 1b”. One possible difference between the value given here and in the original work is normalization. In that case, a ‘Note’ explains how the value was converted. Any additional information, for example, alerting on an unusual configuration or a specific method used in the experiment, is also given as a ‘Note’.
[…]
@@ I.4 What we do and what we do not
After these preliminaries, we will present the spin-qubit figures of merit published until the end of October 2023.
@@ II Coherence times
Figure 1: Typical decay-curve envelopes. (a) Oscillations with a Gaussian decay envelope, typically appearing in Ramsey experiments measuring the inhomogeneous dephasing time $T_{2}^{*}$ in the time domain. (b) Exponential decay typical for relaxation processes, reflecting the relaxation time $T_{1}$. (c) Three curves are plotted for comparison. They are normalized to reach value 1 at the x-axis parameter equal to 0 and value $1/e$ at the x-axis parameter equal to 1. The Gaussian and Lorentzian, the Fourier transforms of the two envelopes given in (a) and (b), arise in probing the decay in the frequency domain. In that case, the horizontal axis on the figure is the frequency $\omega$, see Eq. (II.1).
@@ II.1 Definition and meaning of experimentally-extracted coherence times
In these equations, we Fourier transformed (F.T.) Eqs. (1) and (2) multiplied by $-i$, normalized the results to the same value at zero frequency, and dropped the time-scale subscripts. To discriminate the two cases given in Eq. (II.1) in the frequency domain is even more difficult than in the time domain, since both Gaussian and Lorentzian are quadratic at small frequencies. Compared to the time domain, now the two functions differ more strongly at their tails (that is, for large $\omega$), with algebraic and exponential decay, respectively. However, as already stated, these tails are seldom accessible with the required precision. The three envelope functions discussed so far are plotted in Fig. 1(c) for comparison.
@@ II.2 Measured coherence times of spin qubits
The coherence times of spin qubits are given in Fig. 2. The values had started at around 10 ns inhomogeneous dephasing time in early experiments with qubits in GaAs. Echo techniques can extend the coherence by orders of magnitudes, as can a different material choice. The coherence times published during the year 2023 span five orders of magnitude, depending on the qubit type, material, and protection measures.
Figure 3: Spin coherence times. The panels display the data from Fig. 2 split to the four panels according to the coherence type, denoted in the legend, from the set: inhomogeneous dephasing $T_{2}^{*}$, Hahn echo $T_{2}^{\mathrm{Echo}}$, dynamical decoupling $T_{2}^{\mathrm{DynD}}$, and Rabi decay $T_{2}^{\mathrm{Rabi}}$. In each panel, the horizontal axis uses the qubit type as a discrete category, displacing the points according to their publication date: more recent data are shifted to the right within the non-shaded area which is normalized to the year span 2003-2023. All vertical-axes ranges are the same and the point colors show the material.
To examine the influence of some of these factors, we plotted separately each type of coherence in Fig. 3. The separation allows us to group the values additionally according to the qubit type, being the discrete category on the horizontal axis. To reflect the publication date for an easier comparison, we displace the data within each category horizontally. For example, Fig. 3(a) shows that the most recent electron spin-1/2 qubits implemented in purified silicon reach longer coherence times than singlet-triplet qubits. The coherences of holes have recently got on par with the best ones for singlet-triplet qubits in almost every category. Finally, impurity spins hold record $T_{2}^{*}$ and $T_{2}^{\mathrm{Echo}}$ coherence times.
@@ II.3 Measured coherence times of charge qubits
All types of coherence times of charge qubits, including the relaxation time, are gathered in Fig. 5. In the three panels of the figure, the same data are shown according to the publication date, the device material, and the coherence type, respectively. One can see several differences compared to spin qubits. First, the relaxation times are often comparable to coherence times, unlike for most spin-qubits where they can be pushed to be by far the longest scale. Therefore, the relaxation is a bigger issue for charge qubits even in single-qubit experiments. Second, unlike for spin qubits, the echo techniques do not prolong coherence substantially. Third, there is much less variation among the data: Despite an upward trend in $T_{2}^{*}$ over time, the increase is not dramatic. Since the charge-qubit dephasing time $T_{2}^{*}$ is related to charge noise directly, a clear trend would indicate an overall improvement of the available samples and devices concerning the level of charge noise. Finally, as one would expect based on the nuclear-spin noise being irrelevant for charge qubits, there is no apparent difference between devices made in Si and III-V materials concerning charge-qubit coherence.
@@ III.3 The definition and values of measurement times
We do not review the measurement times of charge qubits because the task of measuring a charge qubit is the task of detecting a charge—typically an elementary charge—in a nanodevice. First, this task is separable from, and not exclusive to, spin qubits, and thus not our focus. Second, the analysis of methods and results of this task is a topic rich enough for a review in its own right.[107, 56, 348] For our purposes, it will be enough to give the following minimum. The charge detection is typically characterized through the detection sensitivity $s$ with a typical number being777 Here are a few numbers that one can find in the literature: $10$ in Ref. [283], $0.07$ in Ref. [6] $10$ in Ref. [359], $30$ in Ref. [90], $63$ in Ref. [57], $0.8$ in Ref. [312], $0.37$ in Ref. [103], $8.2$ in Ref. [387], $4.1$ in Ref. [391], $3.3$ in Ref. [66], $21$ in Ref. [50], $0.60$ in Ref. [298]; all in $10^{-4}\times e/\surd{\mathrm{Hz}}$.
@@ IV.2 Measured values of fidelities
The three panels of Fig. 8 show the published fidelities of the gates, measurements, and initializations, respectively. For the gates, Fig. 8(a), the electron spin-1/2 qubits previously reached the highest fidelities, well above 99.9 %. Using silicon, both natural and isotopically purified, was crucial for this achievement. Recently, these values were overcome and a new record at $9$ has been set by a hole in germanium [194]. There is notable progress in almost every qubit category and increasing the fidelity of single-qubit gates is one of the most impressive achievements within the whole spin-qubit field.
@@ Appendix A Relaxation time notations
the two sets of equations, (A) and (A), correspond to each other.
In the above equations for the density matrix evolution, we dropped the coherent part, which would read $\partial_{t}\mathbf{s}=(g\mu_{B}/\hbar)\mathbf{s}\times\mathbf{B}$, with $g$ being the electron g-factor and $\mathbf{B}$ the magnetic field. Equations (A) are derived in Ref. [83] in a simple model on p. 707-712. The same reference also derives Eq. (A) on p. 759-765, considering a spin qubit coupled to phonons as the decay source, following the standard methods of Ref. [22].
@@ Appendix B Relation of the gate fidelity and the gate quality factor
Let us now consider the coherence decay given by Eq. (A). Assuming, for simplicity, that either the rotation is in the plane of the Bloch sphere perpendicular to the energy-quantization axis, or taking $T_{2}=T_{1}$ and $s_{0z}=0$, the actual density matrix is described by $\mathbf{s}(t)\exp(-t/T_{2})$. The amplitude of the oscillating polarization measured as a function of time will decay exponentially, as $\exp(-t/T_{2})$. The fitted time-scale of the decay is called the “gate decay time” $T_{2}^{\mathrm{Rabi}}$ entering Eq. (21). Evaluating the fidelity with Eq. (14) gives
@@ Appendix C List of tables
Spin coherence Table C LABEL:spinCoherence1-LABEL:spinCoherence4
Charge coherence Table C LABEL:chargeCoherence
Operation times Table C LABEL:operationTime1-LABEL:operationTime2
Operation fidelities Table C LABEL:operationFidelity1-LABEL:operationFidelity2

### v9

@@ I.2 Spin-qubit types, device geometries, material choices
Every value given in this Technical Review (such as a coherence time of 10 ns) is a result of a measurement with a specific device and qubit type, reported in a published reference. [^1] Note that whereas we count arXiv preprints as ‘published references’, the majority of the entries are from peer-reviewed journal publications. The current bibliography for the database contains 14 arXiv preprints, only four of which were uploaded before the year 2024, Refs. [130, 49, 310, 308]. This means, first of all, that we include only numbers explicitly stated and directly measured. We do not include extrapolations to other conditions or materials. Also, we do not usually derive the values ourselves even if it would be possible. For example, if the reference gives the operation time and the dephasing time, but does not state the quality factor, we do not evaluate the latter ourselves. However, if a quantity is discussed and presented as a figure (for example, the operation time implied from oscillations displayed by a resonantly driven qubit), we might include a value read-off from the figure. In such cases, the table entry contains a ‘Note’ with keywords (derived)  or (estimated) , which are explained in Appendix D. With this requirement, every value given in this Technical Review should be easy to find using the reference, given under the attribute ‘Reference’, and within the reference as described by the attribute ‘Source’. An example of the latter is say “page 4 and Fig. 1b”. One possible difference between the value given here and in the original work is normalization. In that case, a ‘Note’ explains how the value was converted. Any additional information, for example, alerting on an unusual configuration or a specific method used in the experiment, is also given as a ‘Note’.
The second group of common characteristics concerns the details of the qubit. We found it useful to categorize the following: the sample material, the geometry of the host, and the qubit type. They belong to the attributes ‘Material’, ‘Host’, and ‘Qubit’, respectively. Whereas the value for material, for example, ‘Si/SiGe’, is self-explanatory, [^2] Qubits based on silicon-oxide structures is one case that needs a comment. A unique identification in this Technical Review for such structures is the value ‘Si/SiO2’ of the attribute ‘Material’. The attribute ‘Host’ is typically also straightforward to assign, being either ‘2D’, for example for an epilayer, or ‘1D’ for finFETs or structures denoted as nanowires by their authors. One can often find further specifications for such devices, such as [complementary-]metal-oxide-semiconductor([C]MOS), silicon-on-insulator(SOI), field-effect-transistor(FET), foundry-compatible, and similar, including their combinations. These specifications hint at the fabrication details and the degree of compatibility with the industrial silicon technology. However, they are sometimes used interchangeably, even within one laboratory. Since the fabrication details are not our focus, we do not include such additional specifications even if given in the original work. in some figures we group several different materials under a common tag, such as ‘Si’. Concerning the host geometry, we discriminate (qubits based on gating) the structures which are (quasi-)‘2D’, for example, a 2D electron gas (2DEG), (quasi-)‘1D’, for example a nanowire, and quasi-zero-dimensional, denoted by ‘imp’, an example being an implanted impurity. Some of these are not clear-cut cases, for example the hut wires with a flat cross-section [398], or some variants of CMOS devices [379]; nevertheless, we assign both to ‘1D’. Finally, perhaps the largest variation exists among the qubit types. We distinguish the charge carrier: conduction electron, valence hole, and atomic impurity; and the spin-encoding: spin-1/2 (‘LD’), singlet-triplet (‘ST’), and hybrid (‘HY’) qubits. We have not found it beneficial to subdivide further the hybrid qubits: everything which is not a spin-1/2 or a singlet-triplet is assigned the value hybrid here. A note might give additional information on the qubit type. The reader would benefit from consulting Appendix D now, to understand the database organization through attributes and values.
@@ I.4 What we do and what we do not
After these preliminaries, we will present the spin-qubit figures of merit published until the end of September 2024.
@@ II Coherence times
Figure 1: Typical decay-curve envelopes. (a) Oscillations with a Gaussian decay envelope, typically appearing in Ramsey experiments measuring the inhomogeneous dephasing time $T_{2}^{*}$ in the time domain. (b) Exponential decay typical for relaxation processes, reflecting the relaxation time $T_{1}$. (c) Three curves are plotted for comparison. They are normalized to reach value 1 at the x-axis parameter equal to 0 and value $1/e$ at the x-axis parameter equal to 1. The Gaussian and Lorentzian, the Fourier transforms of the two envelopes given in (a) and (b), arise in probing the decay in the frequency domain. In that case, the horizontal axis on the figure is the frequency $\omega$, see Eq. (7).
@@ II.1 Definition and meaning of experimentally-extracted coherence times
In these equations, we Fourier transformed (F.T.) Eqs. (1) and (2) multiplied by $-i$, normalized the results to the same value at zero frequency, and dropped the time-scale subscripts. To discriminate the two cases given in Eq. (7) in the frequency domain is even more difficult than in the time domain, since both Gaussian and Lorentzian are quadratic at small frequencies. Compared to the time domain, now the two functions differ more strongly at their tails (that is, for large $\omega$), with algebraic and exponential decay, respectively. However, as already stated, these tails are seldom accessible with the required precision. The three envelope functions discussed so far are plotted in Fig. 1(c) for comparison.
@@ II.2 Measured coherence times of spin qubits
The coherence times of spin qubits are given in Fig. 2. The values had started at around 10 ns inhomogeneous dephasing time in early experiments with qubits in GaAs. Echo techniques can extend the coherence by orders of magnitudes, as can a different material choice. The coherence times published during the year 2024 span five orders of magnitude, depending on the qubit type, material, and protection measures.
Figure 3: Spin coherence times. The panels display the data from Fig. 2 split to the four panels according to the coherence type, denoted in the legend, from the set: inhomogeneous dephasing $T_{2}^{*}$, Hahn echo $T_{2}^{\mathrm{Echo}}$, dynamical decoupling $T_{2}^{\mathrm{DynD}}$, and Rabi decay $T_{2}^{\mathrm{Rabi}}$. In each panel, the horizontal axis uses the qubit type as a discrete category, displacing the points according to their publication date: more recent data are shifted to the right within the non-shaded area which is normalized to the year span 2003-2024. All vertical-axes ranges are the same and the point colors show the material.
To examine the influence of some of these factors, we plotted separately each type of coherence in Fig. 3. The separation allows us to group the values additionally according to the qubit type, being the discrete category on the horizontal axis. To reflect the publication date for an easier comparison, we displace the data within each category horizontally. For example, Fig. 3(a) shows that the most recent electron spin-1/2 qubits implemented in purified silicon reach longer coherence times than singlet-triplet qubits. The coherences of holes have recently got on par with the best ones for singlet-triplet qubits in almost every category. Finally, impurity spins hold record $T_{2}^{*}$, $T_{2}^{\mathrm{Echo}}$, and $T_{2}^{\mathrm{Rabi}}$ coherence times.
@@ II.3 Measured coherence times of charge qubits
All types of coherence times of charge qubits, including the relaxation time, are gathered in Fig. 5. In the three panels of the figure, the same data are shown according to the publication date, the device material, and the coherence type, respectively. One can see several differences compared to spin qubits. First, the relaxation times are often comparable to coherence times, unlike for most spin-qubits where they can be pushed to be by far the longest scale. Therefore, the relaxation is a bigger issue for charge qubits even in single-qubit experiments. Second, unlike for spin qubits, the echo techniques do not prolong coherence substantially. Third, there is much less variation among the data: Despite some upward trend in $T_{2}^{*}$ over time, the increase is not dramatic. Since the charge-qubit dephasing time $T_{2}^{*}$ is related to charge noise directly, a clear trend would indicate an overall improvement of the available samples and devices concerning the level of charge noise. Finally, as one would expect based on the nuclear-spin noise being irrelevant for charge qubits, there is no apparent difference between devices made in Si and III-V materials concerning charge-qubit coherence.
@@ III.3 The definition and values of measurement times
We do not review the measurement times of charge qubits because the task of measuring a charge qubit is the task of detecting a charge—typically an elementary charge—in a nanodevice. First, this task is separable from, and not exclusive to, spin qubits, and thus not our focus. Second, the analysis of methods and results of this task is a topic rich enough for a review in its own right.[116, 59, 378] For our purposes, it will be enough to give the following minimum. The charge detection is typically characterized through the detection sensitivity $s$ with a typical number being777 Here are a few numbers that one can find in the literature: $10$ in Ref. [304], $0.07$ in Ref. [6] $10$ in Ref. [391], $30$ in Ref. [98], $63$ in Ref. [60], $0.8$ in Ref. [338], $0.37$ in Ref. [112], $8.2$ in Ref. [421], $4.1$ in Ref. [426], $3.3$ in Ref. [69], $21$ in Ref. [53], $0.60$ in Ref. [323], $31$ in Ref. [280]; all in $10^{-4}\times e/\surd{\mathrm{Hz}}$.
@@ IV.2 Measured values of fidelities
The three panels of Fig. 8 show the published fidelities of the gates, measurements, and initializations, respectively. For the gates, Fig. 8(a), the electron spin-1/2 qubits previously reached the highest fidelities, well above 99.9 %. Using silicon, both natural and isotopically purified, was crucial for this achievement. Recently, these values were overcome and a new record above $9$ has been set by a hole in germanium [208]. There is notable progress in almost every qubit category and increasing the fidelity of single-qubit gates is one of the most impressive achievements within the whole spin-qubit field.
@@ Appendix A Relaxation time notations
the two sets of equations, (23) and (24), correspond to each other.
In the above equations for the density matrix evolution, we dropped the coherent part, which would read $\partial_{t}\mathbf{s}=(g\mu_{B}/\hbar)\mathbf{s}\times\mathbf{B}$, with $g$ being the electron g-factor and $\mathbf{B}$ the magnetic field. Equations (23) are derived in Ref. [91] in a simple model on p. 707-712. The same reference also derives Eq. (24) on p. 759-765, considering a spin qubit coupled to phonons as the decay source, following the standard methods of Ref. [24].
@@ Appendix B Relation of the gate fidelity and the gate quality factor
Let us now consider the coherence decay given by Eq. (23). Assuming, for simplicity, that either the rotation is in the plane of the Bloch sphere perpendicular to the energy-quantization axis, or taking $T_{2}=T_{1}$ and $s_{0z}=0$, the actual density matrix is described by $\mathbf{s}(t)\exp(-t/T_{2})$. The amplitude of the oscillating polarization measured as a function of time will decay exponentially, as $\exp(-t/T_{2})$. The fitted time-scale of the decay is called the “gate decay time” $T_{2}^{\mathrm{Rabi}}$ entering Eq. (21). Evaluating the fidelity with Eq. (14) gives
@@ Appendix C List of tables
Spin coherence Table C LABEL:spinCoherence1-LABEL:spinCoherence8
Charge coherence Table C LABEL:chargeCoherence1LABEL:chargeCoherence2
Operation times Table C LABEL:operationTime1-LABEL:operationTime4
Operation fidelities Table C LABEL:operationFidelity1-LABEL:operationFidelity5